OpenStudy (anonymous):

Determine the limit of the sequence a_n= n!/ (n+3)!

OpenStudy (blockcolder):

\[\lim_{n \rightarrow \infty} n!/(n+3)!= \lim_{n \rightarrow \infty}1/[(n+1)(n+2)(n+3)]=0.\] :D

OpenStudy (anonymous):

THANKS!! I got that part of the limit, and then my brain stopped.